two remotely controlled unmanned aircraft are launched in opposite directions. One flies east at 78 mph and the other west at 82 mph. How long will it take the aircraft to fly a combined distance of 560 miles.
Answer by stanbon(75887)

two remotely controlled unmanned aircraft are launched in opposite directions. One flies east at 78 mph and the other west at 82 mph. How long will it take the aircraft to fly a combined distance of 560 miles.
-----
They are separating at a rate of (78+82) = 160 mph
------
Ans: time = distanc/rate = 560/160 = 3.5 hours
---------------
Cheers,
Stan H.
================